第46回環論および表現論シンポジウム（2013年）第二報

第 46 回環論および表現論シンポジウムのプログラムが出来ましたのでお知らせいたします。また懇親会もあわせてご案内申し上げます。このシンポジウムは以下の研究費の援助のもとに行われます。

日本学術振興会科学研究費補助金基盤研究 (B) 研究代表者：浅芝秀人 (静岡大学) 日本学術振興会科学研究費補助金基盤研究 (C) 研究代表者：馬場良始 (大阪教育大学) 研究代表者：池畑秀一 (岡山大学) 研究代表者：西中恒和（岡山商科大学）

プログラム責任者菊政 勲 (山口大学)

記

日程 : 2013 年 10 月 12 日 (土)〜14 日 (月・祝)

場所 : 東京理科大学神楽坂キャンパス

〒 162-8601 東京都新宿区神楽坂 1-3

       http://www.tus.ac.jp/info/access/kagcamp.html

・10 月 12 日 (土) 8 号館 3 階 832 教室・10 月 13 日 (日) 2 号館 2 階 221 教室・10 月 14 日 (月・祝) 森戸記念館 第１フォーラム

会場責任者 : 眞田克典 (東京理科大学 sanada@rs.kagu.tus.ac.jp)

プログラム責任者 :菊政 勲 (山口大学 kikumasa@yamaguchi-u.ac.jp)

懇親会 : 日時 : 10 月 13 日 (日) 18:00 〜

会場 : 神楽坂キャンパス８号館学生食堂

会費 : 5,000 円

懇親会参加希望者は 9 月 27 日 (金) までに、 E-mail にて、会場責任者・眞田克典氏 (sanada@rs.kagu.tus.ac.jp) までお知らせ下さい。多数の方のご参加をお待ちしています。

10 月 12 日（土）

9:00–9:30 小原大樹（東京理科大学） One point extension of quiver algebras defined by two cycles and a quantumlike relation

9:30–10:00 小西正秀（名古屋大学） A classification of cyclotomic KLR algebras of type A_{n}^{(1)}

10:15–10:45 古谷貴彦（明海大学）・速水孝夫（北海学園大学） On some finiteness questions about Hochschild cohomology of finite-dimensional algebras

10:45–11:15 佐々木洋城（信州大学） Source algebras and cohomology of block ideals of fintie group algebras

11:30–12:00 伊山修（名古屋大学）

Geigle-Lenzing spaces and canonical algebras in dimension d 12:00–12:30 松田一徳（立教大学・ JST CREST）

Stable set polytopes of trivially perfect graphs

14:00–14:30 加藤希理子（大阪府立大学）・ Peter Jørgensen (Newcastle University)

Triangulated subcategories of Extensions 14:30–15:00 劉裕（名古屋大学）

Hearts of twin cotorsion pairs on exact categories

15:15–15:45 平松直哉（呉工業高専）

Serre subcategories of artinian modules 15:45–16:15 神田遼（名古屋大学）

Specialization orders on atom spectra of Grothendieck categories

16:30–17:20 Manuel Saorín Castaño (Universidad de Murcia)

Resolving subcategories of modules of finite projective dimension over a commutative ring

10 月 13 日（日）

9:00–9:30 平野康之（鳴門教育大学）

On the ring of complex-valued functions on a finite ring 9:30–10:00 松岡学（大阪樟蔭女子大学）

QF rings and direct summand conditions

10:15–10:45 古賀寛尚・星野光男（筑波大学）・亀山統胤（信州大学）

Clifford extensions 10:45–11:15 亀山統胤（信州大学）・星野光男・古賀寛尚（筑波大学）

Group-graded and group-bigraded rings

11:30–12:20 Manuel Saorín Castaño (Universidad de Murcia)

Classical derived functor as fully faithful embeddings

14:00–14:30 竹花靖彦（函館工業高専）

Complements and closed submodules relative to torsion theories 14:30–15:00 源泰幸（大阪府立大学）

Torsion theory and ideals

15:15–15:45 金加喜・松本英鷹・松澤翔（静岡大学）

Defining relations of 3-dimensional quadratic AS-regular algebras 15:45–16:15 秋山諒（静岡大学）

Quantum planes and iterated Ore Extensions

16:30–17:20 Manuel Saorín Castaño (Universidad de Murcia)

The symmetry, period and Calabi-Yau dimension of m-fold mesh algebras

18:00– 懇親会

10 月 14 日（月・祝）

9:00–9:30 相原琢磨（名古屋大学）

Mutation and mutation quivers of symmetric special biserial algebras 9:30–10:00 水野有哉（名古屋大学）

Support tau-tilting modules and preprojective algebras

10:15–10:45 足立崇英（名古屋大学）

Classifying τ-tilting modules over Nakayama algebras 10:45–11:15 加瀬遼一（大阪大学）

On the poset of pre-projective tilting modules over path algebras

11:30–12:00 板垣智洋・眞田克典（東京理科大学） The dimension formula of the cyclic homology of truncated quiver algebras over a field of positive characteristic